OpenStudy (anonymous):

For the given functions f, g and h, find f○g○h and state the exact domain of f○g○h. Please show all of your work. f(x) = 1/x g(x) - lnx h(x) = 2x+15

OpenStudy (anonymous):

h(x) = 2x + 15 g(h(x)) = - ln (2x+15) So what's f(g(h(x)))?

OpenStudy (anonymous):

\[h(x)=2x+15 \text{ domain of } h: \mathbb{R}\] \[g(x)=-\ln(x) \text{ domain of }g =\{x \in \mathbb{R}: x>0\}\] \[g\circ h(x)=-\ln(2x+15) \text{ domain of }g \circ h =\{x \in \mathbb{R}: 2x+15>0\}=\{x \in \mathbb{R}: x>-\frac{15}{2}\}\] \[f(x)=\frac{1}{x} \text{ domain of } f=\{x \in \mathbb{R}: x \neq 0\}\] \[f \circ g \circ h(x)=-\frac{1}{\ln(2x+15)} \text{ domain of } f \circ g \circ h =\{x \in \mathbb{R}: 2x+15>0 \text{ and } \ln(2x+15) \neq 0\}\] \[\ln(2x+15)\neq0 \Rightarrow 2x+15 \neq 1 \Rightarrow x \neq -7\] \[\text{ domain of } f \circ g \circ h =\{x \in \mathbb{R}: x \in ]-\frac{15}{2},-7[ \cup]-7,+\infty[ \}\]

OpenStudy (anonymous):

vood is this the answer?